OpenStudy (anonymous):

Jared has a wooden board that is 11 ft long. He wants to cut it into two pieces so that the longer piece is 4 ft less than 4 times the length of the shorter piece. How long will each piece of board be? Shorter piece: ft Longer piece: ft

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@wmj259

OpenStudy (wmj259):

Ooh a toughie

OpenStudy (anonymous):

uh oh lol a toughie? your mean a tough question? xD toughie tough guy both together mean toughie or looks like touhgie xD im so mean

OpenStudy (anonymous):

you not your mean xD didnt mean that to happen xD

OpenStudy (anonymous):

any idea?

OpenStudy (wmj259):

I'm trying to make it into an equation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think i have had this one before this first one is like 8 and the second one is like -32, -36, or -39 one of those three

OpenStudy (wmj259):

You cannot have a negative length.

OpenStudy (mathstudent55):

The short piece is x. The long piece is "4 ft less than 4 times the length of the shorter piece." Using x for the short piece, can you write an expression to show the length of the longer piece?

OpenStudy (anonymous):

o xD

OpenStudy (mathstudent55):

4 ft less than 4 times x 4 times x is 4x 4 ft less than 4x is 4x - 4 Ok?

OpenStudy (wmj259):

@mathstudent55 I thought we can have 11=2T then T=X+Y then Y=4X-4 does that sound right?

OpenStudy (anonymous):

ok

OpenStudy (wmj259):

@mathstudent55 congrats

OpenStudy (mathstudent55):

11 = 2T will not work because that means each T is equal. You can use some of what you wrote: Let the pieces be x and y. x + y = 11 y = 4x - 4

OpenStudy (mathstudent55):

Now take y (which is 4x - 4) and substitute it into the y of x + y = 11. You get: x + (4x - 4) = 11 x + 4x - 4 = 11 5x - 4 = 11 5x = 15 x = 3 The short piece is 3. The long piece has to be 8 since 3 + 8 = 11

OpenStudy (wmj259):

Oh I assumed each piece would be the same. So 11=2T is not right.

OpenStudy (mathstudent55):

Now let's make sure we are correct. The longer piece is 4 ft less than 4 times the short piece. 4 * 3 ft = 12 ft 12 ft - 4 ft = 8 ft It works out. The lengths are: 3 ft and 8 ft

OpenStudy (wmj259):

@mathstudent55 good job.

OpenStudy (mathstudent55):

You are correct now. 11 = 2T is not correct. The rest of the equations you wrote were good. In fact, I used them to solve the problem.

OpenStudy (mathstudent55):

Thanks, you too!

OpenStudy (anonymous):

yay i knew there was an 8 somewhere ^_^ im so cool xD im so lame accually

OpenStudy (wmj259):

*facepalm* ;)